OpenStudy (anonymous):

find dy/dx y=integral from cubic root of x TO pi/3 OF cost^3 help!!

OpenStudy (anonymous):

dt

OpenStudy (lgbasallote):

\[\large \int_{\sqrt[3]x}^{\pi/3} \cos(t^3)dt\] is this the question??

OpenStudy (anonymous):

yes @lgbasallote

OpenStudy (lgbasallote):

are you sure it's not \[\large \int_{\sqrt[3]x} ^{\pi/3} (\cos t)^3 dt\]

OpenStudy (anonymous):

nooooo, the 1st one

OpenStudy (anonymous):

u there?

OpenStudy (anonymous):

???